40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Nov 19 14:59:48 2018, Chub, Update, VAC, Vent 81 
    Reply  Mon Nov 19 17:09:00 2018, Steve, Update, VAC, Vent 81 beforeVent82.pngvent81completed.png
       Reply  Mon Nov 19 22:01:50 2018, gautam, Update, VAC, Loose nut on valve IMG_7171.JPG
          Reply  Tue Nov 20 13:13:01 2018, gautam, Update, VAC, IMC alignment is okay 
       Reply  Mon Nov 26 15:58:48 2018, Steve, Update, VAC, Vent 81 
          Reply  Mon Nov 26 17:16:27 2018, gautam, Update, SUS, EY chamber work IMG_5251.JPG
          Reply  Thu Dec 20 10:28:43 2018, gautam, Update, VAC, Heavy doors back on for pumpdown 82 
Message ID: 14319     Entry time: Mon Nov 26 17:16:27 2018     In reply to: 14318
Author: gautam 
Type: Update 
Category: SUS 
Subject: EY chamber work 

[steve, rana, gautam]

  • PSL and EY 1064nm laser (physical) shutters on the head were closed so that we and sundance crew could work without laser safety goggles. EY oplev laser was also turned off.
  • Cylindrical heater setup removed:
    • heater wiring meant the heater itself couldn't be easily removed from the chamber
    • two lenses and Al foil cylinder removed from chamber, now placed on the mini-cleanroom table.
  • Parabolic heater is untouched for now. We can re-insert it once the test mass is back in, so that we can be better informed about the clipping situation.
  • ETMY removed from chamber.
    • EQ stops were engaged.
    • Pictures were taken
    • OSEMs were removed from cage, placed in foil holders.
    • Cage clamps were removed after checking that marker clamps were in place.
    • Optic was moved first to NW corner of table, then out of the vacuum onto the mini-cleanroom desk Chub and I had setup last week.
    • Hoepfully there isn't an earthquake. EY has been marked as off-limits to avoid accidental bumping / catasrophic wire/magnet/optic breaking.
    • We sealed up the mini cleanroom with tape. F.C. cleaning tomorrow or at another opportune moment.
    • Light door was put back on for the evening.

Rana pointed out that the OSEM cabling, because of lack of a plastic shielding, is grounded directly to the table on which it is resting. A glass baking dish at the base of the seismic stack prevents electrical shorting to the chamber. However, there are some LEMO/BNC cables as well on the east side of the stack, whose BNC ends are just lying on the base of the stack. We should use this opportunity to think about whether anything needs to be done / what the influence of this kind of grounding is (if any) on actuator noise.

Steve also pointed out that we should replace the rubber pads which the vacuum chamber is resting on (Attachment #1, not from this vent, but just to indicate what's what). These serve the purpose of relieving small amounts of strain the chamber may experience relative to the beam tube, thus helping preserve the vacuum joints b/w chamber and tube. But after (~20?) years of being under compression, Steve thinks that the rubber no longer has any elasticity, and so should be replaced.

Attachment 1: IMG_5251.JPG  5.099 MB  Uploaded Mon Dec 3 16:23:52 2018  | Hide | Hide all
IMG_5251.JPG
ELOG V3.1.3-